ULTRAFAST STRUCTURAL DYNAMICS OF 1,3-CYCLOHEXADIENE: ELECTRONIC STATE DEPEN-DENCE

\underline{CHRISTINE C. BÜHLER}, MICHAEL P. MINITTI, SANGHAMITRA DEB, PETER M. WEBER, Department of Chemistry, Brown University, Providence, Rhode Island 02912; JIE BAO, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

The ultrafast structural dynamics of 1,3-cyclohexadiene has been investigated using structurally sensitive Rydberg electron binding energies. Excitation to the 1B state and the 3p Rydberg state yielded different structural responses. In both experiments, the structural dynamics of the molecular core are reflected by time-dependent shifts of the Rydberg electron binding energy. Structural distortions associated with 3p-excitation cause a dynamical shift in the p_{x}- and p_{y}-binding energies by 8 and 25 meV/ps respectively, whereas after excitation into 1B more severe structural transformations along the ring-opening coordinate produce binding energy shifts at a rate of 65 meV/ps.